OpenStudy (calculusxy):

A shelf at a bookstore displays 27 books. Of these 27 books, 9 of the books are nonfiction books. The store owner adds 6 new fiction books to the shelf and wants the ratio of fiction to nonfiction books to remain the same. What is the total number of nonfiction books that the store should display, after the new fiction books are added?

OpenStudy (calculusxy):

I have a feeling that it is 18.

OpenStudy (anonymous):

I think you'll need to use variables.

OpenStudy (calculusxy):

Well I thought that nine goes to twenty seven three times, and so six will have to be multiplied by 3 to get 18.

OpenStudy (anonymous):

n for nonfiction books. f for fiction books.

OpenStudy (calculusxy):

okay so first of all, is 18 the wrong answer?

OpenStudy (anonymous):

Hold on one second, I am going to solve it on my own and then try to explain.

OpenStudy (anonymous):

The initial ratio is: \[\frac{9}{27} \]

OpenStudy (anonymous):

When we add the \(6\) fiction books, then we have: \[ \frac{9}{27+6} \]

OpenStudy (anonymous):

We need to add a certain number of non-fiction books, we'll suppose that we need to add \(n\) of them. Then we want: \[ \frac{9}{27} = \frac{9+n}{27+6+n} \]

OpenStudy (calculusxy):

How would I solve for n?

OpenStudy (anonymous):

One thing you can do is cross multiply.

OpenStudy (anonymous):

Basically, \[ 9\times (27+6+n) = 27\times (9+n) \]

OpenStudy (calculusxy):

So 9(27+6+n) and 27(9+n)?

OpenStudy (calculusxy):

And I need to use the distributive property right?

OpenStudy (anonymous):

yes

OpenStudy (calculusxy):

I got 243 + 54 + 9n and 243 + 27n. Am I Correct?

OpenStudy (anonymous):

You still need to solve for \(n\).

OpenStudy (calculusxy):

Okay. 243 + 54 + 9n = 243 + 27n

OpenStudy (calculusxy):

I subtracted 9n from both sides. Got: 243 + 54 = 243 + 18n Subtracted 243 from both sides, got 54= 18n Divided n from both sides, and n equals to 3

OpenStudy (calculusxy):

So is it 3?

OpenStudy (anonymous):

Well, does \[ \frac{9}{27} = \frac{9+3}{27+6+3} \]?